# Cream: Consistency Regularized Self-Rewarding Language Models

### v2

@@ 4.3.2 Reliability of Self-Consistency
The most straightforward way to enhance the rewarding and ranking accuracy is by incorporating external reward models, such as the SRLM variant “Oracle” used in our experiments. Recall the theoretical analysis in Eq. (3.8) suggests that we can mitigate the rewarding bias issue by calculating the ranking consistency between the current model and other available baseline reward models (BRMs). However, it is not feasible to have access to external reward models acting as BRMs in the self-rewarding scenario. Thus, we instead propose to use the last iteration’s model as the BRM to measure the rewarding consistency.
Choice of Baseline Reward Model. To measure the impacts on CREAM of using different BRMs, we fine-tune the M1 model using CREAM with two different BRMs: the rewarding function of Oracle and the model from the last iteration (M0). As shown in Table 3, using a strong reward model as the BRM can bring a better regularization effect, especially for Llama2. However, we find that the last iteration’s model also provides a reasonably reliable consistency signal for Llama3. We attribute this to Llama3’s inherently stronger foundational alignment and better internal consistency, which allows it to effectively utilize itself without needing an external reward model.
@@ 4.3.3 Consistency Measurement
Table 4: Performance of CREAM using different consistency measurements with Llama3.
@@
Besides the adopted Kendall $\tau$ coefficient, other metrics can also be used to measure the consistency between two preference ranking lists, such as Spearman coefficient (Spearman, 1904) and the aforementioned TopOrder method. Table 4 compares the performance of CREAM using different consistency measurements. We can observe that: (1) Compared to M1, all these measurements are effective with CREAM, indicating the generalization and applicability of our regularized training. (2) Kendall correlation coefficient generally yields higher scores across various datasets compared to Spearman and TopOrder measurement methods. (3) The differences in performance highlight the sensitivity of these consistency measurements. Specifically, the Spearman coefficient appears slightly less robust than Kendall’s $\tau$, as analyzed in Croux & Dehon (2010). Meanwhile, TopOrder focuses only on top-1 and bottom-1 rankings, limiting its evaluation scope.
4.3.4 Case Study
Table 5: Case study of the generated response sampled from SRLM and CREAM models.
Question Today is the 5th of March, what is the date in 6 weeks?
Method Response sampled from M2 Ranking
[…]
Table 5 presents a case study to show the quality of the generated responses and the consistency of rankings across iterations. We find that the quality of top-ranked response generated by CREAM is significantly better than that of SRLM, showing more accurate and concise reasoning. It also shows that CREAM provides consistent rankings between M2 and M3, while SRLM exhibits inconsistent rankings, possibly due to M3 being fine-tuned on noisy preference data annotated by M2.
@@ 5 Conclusion
In this paper, we first formulate a generalized iterative preference fine-tuning framework for self-rewarding language models (SRLMs). Then, we highlight the rewarding bias that emerges from overconfident preference labeling, which is particularly problematic for 7B level smaller LLMs. This rewarding bias results in the accumulation of noisy and unreliable preference data, harming the preference training and hindering alignment performance of LLMs. To address this issue, we proposed the CREAM, which leverages the rewarding consistency across different iterations as a regularization signal. This approach allows the model to learn more selectively, emphasizing reliable preference data and avoiding overconfidence in preference labeling. Experimental results on five natural language benchmark datasets demonstrate the effectiveness of the proposed method in mitigating the rewarding bias issue of SRLMs and improving the alignment performance of LLMs.
Appendix A Proof of the theorems and lemmas
A.1 Proof of Theorem 3.1
@@ Appendix A Proof of the theorems and lemmas
$\mathcal{L}(\bm{\theta}_{t},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t}).$ (A.1)
@@
$\mathcal{L}(\bm{\theta}_{t+1},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t+1}).$ (A.2)
Connecting Eq. (A.1) with Eq. (A.2) yields that the loss function $\mathcal{L}(\bm{\theta},z)$ is monotonically decreasing, i.e.
$\cdots\leq\mathcal{L}(\bm{\theta}_{t+1},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t})\leq\cdots.$ (A.3)
@@
$\quad+\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}P_{\bm{\theta}_{t}}(z=1)\left[\log P_{\bm{\theta}}(z=1)+\log P_{\bm{\theta}}(z=0)\right],$ (A.4)
where the second equation decompose the expectation $\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}$ into two expectation $\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}$ + $\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}$, the third equation extract the event $\mathbf{y}\geq\mathbf{y}^{\prime}$ as distribution $P_{\bm{\theta}_{t}}(z)$. Then Eq. (A.4) can be further written by
@@
$\quad+P_{\bm{\theta}_{t}}(z=0)\left[\log P_{\bm{\theta}}(z=1)+\log P_{\bm{\theta}}(z=0)\right],$ (A.5)
since both $\mathbf{y},\mathbf{y}^{\prime}$ are generated from $\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$, $P_{\bm{\theta}_{t}}(z=0)=P_{\bm{\theta}_{t}}(z=1)=0.5$. Thus Eq. (A.5) becomes
$\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\mathcal{L}_{\text{reg}}(\bm{\theta};\mathbf{y},\mathbf{y}^{\prime},\mathbf{x})=2\mathrm{KL}(P_{\bm{\theta}}(z)\parallel P_{\bm{\theta}_{t}}(z))=2\mathrm{KL}(P_{\bm{\theta}}(z)\parallel u(z)),$ (A.6)
@@
$\operatorname{\mathds{1}}[(J_{ij}-J_{i^{\prime}j})(K_{ij}-K_{i^{\prime}j})>0]$ $=\operatorname{\mathds{1}}[(r_{ij}-r_{i^{\prime}j})(r^{\prime}_{ij}-r^{\prime}_{i^{\prime}j})>0]$ (A.7)
$\operatorname{\mathds{1}}[(J_{ij}-J_{i^{\prime}j})(K_{ij}-K_{i^{\prime}j})<0]$ $=\operatorname{\mathds{1}}[(r_{ij}-r_{ij^{\prime}})(r^{\prime}_{ij}-r^{\prime}_{i^{\prime}j})<0].$ (A.8)
Since $r_{ij}>r_{i^{\prime}j}$ yields $\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}$ under the input prompt $xb_{j}$ and language model $\bm{\theta}_{t}$, Eq. (A.7) becomes
@@
$\quad+\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}].$ (A.9)
@@
$=0+2\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}](1-2\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}])]$ (A.10)
where the second equation merge the terms together, and the third equation is due to the fact $\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}]+\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}]=1$, the forth equation reorganize the term and the fifth equation is due to the fact that $\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]-\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]]=0$ due to symmetry. Similarly by reversing the $\prec$ and $\succ$, we can write Eq. (A.10) by
@@
$=0+2\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}](1-2\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}])].$ (A.11)
Adding Eq. (A.10) and Eq. (A.11) together yields
@@
$=2-8\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}]],$ (A.12)

### v3

@@ 4.3.2 Reliability of Self-Consistency
The most straightforward way to enhance the rewarding and ranking accuracy is by incorporating external reward models, such as the SRLM variant “Oracle” used in our experiments. The theoretical analysis in Eq. (3.8) suggests that we can mitigate the rewarding bias issue by calculating the ranking consistency between current model and other available baseline reward models (BRMs). However, it is not always feasible to have access to such external reward models in practice, such as the self-rewarding scenario. Thus, we instead propose to use the last iteration’s model as the BRM to measure the consistency of rewarding.
Choice of Baseline Reward Model. To measure the impacts on CREAM of using different BRMs, we fine-tune the M1 model using CREAM with two different BRMs: the rewarding function of Oracle and the model from the last iteration (M0). As shown in Table 3, using a strong reward model as the BRM can bring better regularization effect, especially for Llama-2. However, we find that the last iteration’s model also provides a reasonably reliable consistency signal for Llama-3. We attribute this to Llama-3’s inherently stronger foundational alignment and better internal consistency, which allows it to effectively utilize itself without needing an external reward model.
@@ 4.3.3 Consistency Measurement
Table 4: Performance of CREAM using different consistency measurements with Llama-3.
@@
Besides the adopted Kendall $\tau$ coefficient, other metrics can also be used to measure the consistency between two preference ranking lists, such as Spearman coefficient (Spearman, 1904) and the aforementioned TopOrder method. We conduct a comparison experiments of using different consistency measurement methods in Table 4. We can observe that: (1) All these measurements are effective with CREAM, indicating the generalization and applicability of our regularized training approach. (2) Kendall correlation coefficient generally yields higher scores across various datasets compared to Spearman and TopOrder methods. (3) The differences in performance highlight the sensitivity of these consistency measurements. Specifically, the Spearman coefficient appears slightly less robust than Kendall’s $\tau$, as analyzed in Croux & Dehon (2010). Meanwhile, TopOrder focuses only on top-1 and bottom-1 rankings, limiting its evaluation scope.
@@ 5 Conclusion
In this paper, we first formulate a generalized iterative preference fine-tuning framework for self-rewarding language models (SRLMs), which can also be applied to other iterative preference training methods. Then, we highlight the rewarding bias that arises from overconfident preference labeling, which is particularly problematic for small LLMs, such as those with 7B parameters. This rewarding bias results in the accumulation of noisy and unreliable preference data, harming the preference training and hindering alignment performance of LLMs. To address this issue, we proposed the Consistency Regularized sElf-Rewarding lAnguage Model (CREAM), which leverages the consistency of rewarding across different iterations as a regularization signal. This approach allows the model to learn more selectively, emphasizing reliable preference data and avoiding overconfidence in preference labeling. Extensive experimental results and analysis on various natural language benchmark datasets demonstrate the effectiveness of the proposed method in mitigating the rewarding bias issue and improving the alignment performance of SRLMs. We believe that these findings can provide valuable insights for future research on self-improvement methods of LLM alignment.
Appendix A Limitations
In this section, we discuss the limitations of our work and propose potential solutions for future research. First, our method requires full fine-tuning of the model over multiple iterations, which is both time-consuming and computationally intensive. For future research, we plan to explore Parameter-Efficient Fine-Tuning (PEFT) methods to reduce training costs. Second, our approach primarily focuses on small (7B level) language models, which may limit its applicability. However, the proposed method is designed to be general and can be applied to models of any size. Third, it is worth investigating more complex consistency-regularized self-rewarding scenarios, such as how to assign different weights to various models during the self-rewarding process based on their importance, and enable these models to collaborate to improve the rewarding accuracy.
Appendix B Proof of the theorems and lemmas
B.1 Proof of Theorem 3.1
@@ Appendix B Proof of the theorems and lemmas
$\mathcal{L}(\bm{\theta}_{t},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t}).$ (B.1)
@@
$\mathcal{L}(\bm{\theta}_{t+1},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t+1}).$ (B.2)
Connecting Eq. (B.1) with Eq. (B.2) yields that the loss function $\mathcal{L}(\bm{\theta},z)$ is monotonically decreasing, i.e.
$\cdots\leq\mathcal{L}(\bm{\theta}_{t+1},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t+1})\leq\mathcal{L}(\bm{\theta}_{t},z_{t})\leq\cdots.$ (B.3)
@@
$\quad+\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}P_{\bm{\theta}_{t}}(z=1)\left[\log P_{\bm{\theta}}(z=1)+\log P_{\bm{\theta}}(z=0)\right],$ (B.4)
where the second equation decompose the expectation $\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}$ into two expectation $\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}$ + $\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}$, the third equation extract the event $\mathbf{y}\geq\mathbf{y}^{\prime}$ as distribution $P_{\bm{\theta}_{t}}(z)$. Then Eq. (B.4) can be further written by
@@
$\quad+P_{\bm{\theta}_{t}}(z=0)\left[\log P_{\bm{\theta}}(z=1)+\log P_{\bm{\theta}}(z=0)\right],$ (B.5)
since both $\mathbf{y},\mathbf{y}^{\prime}$ are generated from $\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$, $P_{\bm{\theta}_{t}}(z=0)=P_{\bm{\theta}_{t}}(z=1)=0.5$. Thus Eq. (B.5) becomes
$\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\mathcal{L}_{\text{reg}}(\bm{\theta};\mathbf{y},\mathbf{y}^{\prime},\mathbf{x})=2\mathrm{KL}(P_{\bm{\theta}}(z)\parallel P_{\bm{\theta}_{t}}(z))=2\mathrm{KL}(P_{\bm{\theta}}(z)\parallel u(z)),$ (B.6)
@@
$\operatorname{\mathds{1}}[(J_{ij}-J_{i^{\prime}j})(K_{ij}-K_{i^{\prime}j})>0]$ $=\operatorname{\mathds{1}}[(r_{ij}-r_{i^{\prime}j})(r^{\prime}_{ij}-r^{\prime}_{i^{\prime}j})>0]$ (B.7)
$\operatorname{\mathds{1}}[(J_{ij}-J_{i^{\prime}j})(K_{ij}-K_{i^{\prime}j})<0]$ $=\operatorname{\mathds{1}}[(r_{ij}-r_{ij^{\prime}})(r^{\prime}_{ij}-r^{\prime}_{i^{\prime}j})<0].$ (B.8)
Since $r_{ij}>r_{i^{\prime}j}$ yields $\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}$ under the input prompt $xb_{j}$ and language model $\bm{\theta}_{t}$, Eq. (B.7) becomes
@@
$\quad+\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}].$ (B.9)
@@
$=0+2\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}](1-2\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}])]$ (B.10)
where the second equation merge the terms together, and the third equation is due to the fact $\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}]+\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}]=1$, the forth equation reorganize the term and the fifth equation is due to the fact that $\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]-\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]]=0$ due to symmetry. Similarly by reversing the $\prec$ and $\succ$, we can write Eq. (B.10) by
@@
$=0+2\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}](1-2\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}])].$ (B.11)
Adding Eq. (B.10) and Eq. (B.11) together yields
@@
$=2-8\mathbb{E}_{\mathbf{y}_{ij},\mathbf{y}_{i^{\prime}j}\sim\pi_{\bm{\theta}_{t}}(\cdot|\mathbf{x}_{j})}[\operatorname{\mathds{1}}[\mathbf{y}_{ij}\succ\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t}]\operatorname{\mathds{1}}[\mathbf{y}_{ij}\prec\mathbf{y}_{i^{\prime}j}|\mathbf{x}_{j},\bm{\theta}_{t-1}]],$ (B.12)